### v1

# ASVD: Activation-aware Singular Value Decomposition for Compressing Large Language Models

Abstract
This paper explores a new post-hoc training-free compression paradigm for compressing Large Language Models (LLMs) to facilitate their wider adoption in various computing environments. We delve into the challenges of LLM compression, notably their dependency on extensive training data and computational resources. We propose a training-free approach dubbed Activation-aware Singular Value Decomposition (ASVD) to address these limitations. ASVD effectively manages activation outliers by adjusting the weight matrix based on the activation distribution, improving decomposition accuracy and efficiency. Our method also addresses the varying sensitivity of different LLM layers to decomposition, with an iterative calibration process for optimal layer-specific decomposition. Experiments demonstrate that ASVD can compress network by 10%-20% without losing reasoning capacities. Additionally, it can be seamlessly integrated with other LLM compression paradigms, showcasing its flexible compatibility. Code and compressed models are available at ASVD4LLM.
1 Introduction
Figure 1: Our post-training LLM decomposition method is orthogonal to existing LLM compression techniques, enabling it to function as a versatile and plug-and-play solution for prevalent compression paradigms, including popular quantization methods.
Numerous strategies have been proposed to mitigate the memory consumption issues in Large Language Models (LLMs), as surveyed by Zhu et al. [2023]. These methods predominantly fall into two categories: neural network compression and system optimizations. In terms of neural network compression, techniques such as weight quantization [Dettmers et al., 2022, Shang et al., 2023], network pruning [Frantar & Alistarh, 2023], and knowledge distillation [Agarwal et al., 2023] have been extensively explored. Meanwhile, system optimizations focus on efficient memory management, with cache mechanisms like FlashAttention [Dao et al., 2022] and PagedAttention [Kwon et al., 2023]. Distinct from these approaches, the paradigm of low-rank decomposition is less explored but promising. This technique involves approximating the weight matrices in neural networks with matrices of lower rank, effectively reducing the model size. Given the massive number of parameters in LLMs, low-rank decomposition offers significant potential for memory reduction. Despite its potential, it remains a relatively underutilized approach in the field of LLM compression. Furthermore, low-rank decomposition is complementary to existing LLM compression techniques by further compressing quantized or pruned models, enhancing overall efficiency [Cheng et al., 2017, Deng et al., 2020].
From the perspective of network compression, traditional low-rank decomposition methods typically adhere to a straightforward process: initially training the original model and subsequently fine-tuning the decomposed model [Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021, Hsu et al., 2022]. While this approach is effective, it is resource-intensive and requires the entire training dataset and substantial computational power for end-to-end backpropagation. Applying this method to Large Language Models (LLMs) would encounter major challenges. Firstly, the training data for LLMs may not always be readily available, often restricted by privacy and commercial considerations. Secondly, the training process for these models is notoriously expensive, both in terms of time and computational resources. Given these constraints, the concept of “training-free” compression emerges as a more viable approach for LLMs [Zhu et al., 2023]. This approach includes methods like LLM post-training quantization[Dettmers et al., 2022, Yuan et al., 2023] and LLM post-training pruning [Frantar & Alistarh, 2023], which compress LLMs without the need for extensive retraining. These training-free methods offer a more practical solution for efficiently compressing LLMs.
To realize low-rank post-training decomposition in LLMs, we conduct an extensive analysis of the baseline methods for LLM decomposition. We first observe that straightforward application of existing low-rank decomposition techniques, which typically necessitate training, proves ineffective for LLMs [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015, Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021]. Digging into the failures, we reveal several challenges unique to post-training decomposition for LLMs. One key challenge involves managing outliers in the activations, which significantly intensifies the decomposition error. This challenge mirrors findings in LLM quantization research [Lin et al., 2023, Kim et al., 2023], highlighting the importance of handling such outliers in LLMs. To address these challenges, we introduce Activation-aware Singular Value Decomposition (ASVD). Specifically, ASVD incorporates the distribution of activations into the decomposition process. It achieves this by scaling the values in the weight matrix column-wisely, based on the distribution patterns observed across input and output channels. This adjustment proves particularly beneficial for channels with outlier activations, allowing the ASVD to allocate enhanced focus to these specific weights. Such targeted adjustment facilitates a more accurate reconstruction of the weight matrix, thereby minimizing prediction errors and improving the overall efficiency of the decomposition process.
Our research further delves into the varying sensitivity of different LLM layers to decomposition. We discovered that weights in Multi-Head Attention (MHA) layers [Vaswani et al., 2017] tend to be more resilient to decomposition compared to those in Multi-Layer Perceptron (MLP) layers. This sensitivity is not uniform across all layers, prompting us to develop an method to assign the compression ratio for each layer. This method assesses each layer’s sensitivity to decomposition at different ranks, enabling us to assign a suitable rank for optimal decomposition. This process is notably efficient, requiring only a limited sample set for evaluation. Additionally, we optimize inference by absorbing singular values after decomposition, resulting in computationally efficient matrix operations. Our approach evenly distributes singular values into other U and V matrices, benefiting weight quantization by reducing quantization errors.
We applied our method to open-access LLMs, specifically LLaMA [Touvron et al., 2023a] and LLaMA-2 [Touvron et al., 2023b]. Our experiments reveal that ASVDcan effectively compress network weights by approximately 10%-20%, with only a marginal 1% accuracy loss on benchmarks like MMLU [Hendrycks et al., 2020]. This compression is achieved in a training-free manner, exemplified by the decomposition process for LLaMA-7b, which takes merely 4 hours on an Nvidia A6000 GPU. Importantly, we conducted preliminary validations of ASVD compatibility with 4/6/8-bit quantization. ASVD emerges as an orthogonal paradigm for LLM compression, seamlessly integrating as a plug-and-play technique that aligns with existing LLM quantization methods.
2 Related Work
2.1 Large Language Model Compression
The field of model compression for Large Language Models (LLMs) has seen a surge of innovative techniques aimed at mitigating the substantial computation and memory requirements these models demand [Zhu et al., 2023]. Various methods have emerged to address this challenge, each taking a unique approach to reduce the memory footprint of LLMs. These methods primarily fall into three categories: weight quantization [Courbariaux et al., 2015, Dettmers et al., 2022], network pruning [LeCun et al., 1989, Frantar & Alistarh, 2023], and knowledge distillation [Hinton et al., 2015, Agarwal et al., 2023]. For the wide body of research on LLM compression, please refer to [Zhu et al., 2023] for the comprehensive survey. Among these methods, weight quantization has gained significant traction in the context of LLMs due to its effectiveness. However, despite its popularity as a neural network compression technique, low-rank factorization has not been extensively explored in the realm of LLMs. Recognizing this gap, we introduce a novel low-rank decomposition method tailored specifically for decomposing the weight matrices of LLMs in a training-free manner. This approach aims to provide an efficient and effective solution for LLM compression, leveraging the untapped potential of low-rank factorization in this domain.
2.2 Low-rank Decomposition
The use of low-rank matrix decomposition as a technique for compressing Deep Neural Networks (DNNs) represents a straightforward yet effective approach, garnering considerable attention within both scientific computing and machine learning domains. In recent years, the challenge of efficiently compressing and accelerating large-scale neural networks via low-rank methods has become a focal point of research. This has led to significant advancements in developing and refining low-rank factorization strategies tailored for DNNs [Schotthöfer et al., 2022].
In the realm of low-rank decomposition for neural network compression, existing methods can be broadly classified into two categories: fixed low rank and variable low rank approaches. Fixed rank methods typically involve decomposing weight matrices of pre-trained networks using techniques like Singular Value Decomposition (SVD) or tensor decomposition, followed by fine-tuning the factorized network [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015]. They also involve constraining weight matrices to maintain a fixed low rank during training [Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021], or constructing layers as linear combinations of layers with varying ranks [Ioannou et al., 2015]. A notable limitation of these methods is the introduction of matrix decomposition rank as an additional hyperparameter requiring fine-tuning. In contrast, rank-adaptive methods address this limitation by automatically determining and adjusting the low-rank structure. In particular, Kim et al. [2015, 2019] apply heuristics search to pre-determine the decomposition rank, while Wen et al. [2017] learn low-rank weights through a loss function penalizing approximated matrix ranks. These approaches remove the need for selecting hyper-parameters of LLMs and highlight the potential of low-rank decomposition as an effective tool for model compression, paving the way towards more efficient and scalable Large Language Models (LLMs).
In this work, we introduce our ASVD, a post-training LLM decomposition approach enabling the adaptive determination of SVD ranks to optimize the matrix approximations based on feature activations.
Difference with TensorGPT [Xu et al., 2023]. In the content of LLM compression via decomposition, the most related work is the concurrent TensorGPT [Xu et al., 2023, Zhu et al., 2023], in which the embedding layer of LLMs is compressed through Tensor-Train Decomposition (TTD) [Oseledets, 2011] in order to store large embeddings in a low-rank tensor format, with much fewer parameters. However, there are several differences between those two methods: (1) Unlike TensorGPT which focuses solely on the token embedding matrix, ASVDaims to compress the entire weight spectrum of LLMs. This holistic approach addresses a more critical aspect of LLM compression, as highlighted in recent studies [Lin et al., 2023, Kim et al., 2023]; (2) From the perspective of low-rank decomposition categorization, our method can realize the low-rank decomposition in a rank-adaptive manner, contrasting with the fixed or predetermined ranks used in TensorGPT.
3 Method
In this section, we initially outline the general process of implementing Singular Value Decomposition (SVD) for weight matrix compression in LLMs, as detailed in Section 3.1. Following this, Section 3.2 identifies and discusses two major challenges associated with implementing SVD-based methods in LLMs. To address these challenges, we propose our Activation-aware Singular Value Decomposition (ASVD) and Sensitivity-aware Truncation Rank Searching (STRC) in Sections 3.3 and 3.4, respectively.
3.1 Naïve SVD for Compressing Weight Matrix
A large language model (LLM) is primarily based on the Transformer model [Vaswani et al., 2017]. A significant amount of memory and computational overhead is incurred by Transformers’ linear weight matrices, especially in its multi-head attention (MHA) and feed-forward layers [Tay et al., 2022]. Consequently, we focus primarily on decomposing these linear weight matrices in this paper. Singular Value Decomposition (SVD) is employed for this purpose, which involves decomposing a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ into three matrices: $\mathbf{U}$, $\mathbf{\Sigma}$, and $\mathbf{V}^{T}$, such that $\mathbf{W}\approx\mathbf{W}\mathbf{\Sigma}\mathbf{V}^{T}$), where $\mathbf{\Sigma}$ is an $m\times n$ diagonal matrix, the diagonal values in $\Sigma$ are the singular values of $\mathbf{W}$, and $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are corresponding right and left singular vector matrices, respectively [Demmel, 1997].
The SVD compression process for a weight matrix can be summarized in three steps: Decomposition: Factorize $\mathbf{W}$ using SVD. Truncation: Retain the top $k$ singular values and their corresponding right and left singular vectors. This results in approximated matrices $\mathbf{U}_{k}$, $\mathbf{\Sigma}_{k}$, and $\mathbf{V}_{k}^{T}$, where the right singular vector matrix $\mathbf{U}_{k}$ is $m\times k$, singular $\mathbf{\Sigma}_{k}$ is $k\times k$, and left singular vector matrix $\mathbf{V}_{k}^{T}$ is $k\times n$. The choice of $k$ is critical in balancing the compression ratio and the compressed model’s performance. Reconstruction: Reconstruct an approximated weight matrix:
$\mathbf{W}_{k}=\mathbf{U}_{k}\mathbf{\Sigma}_{k}\mathbf{U}_{k}^{T}.$ (1)
The primary factor impacting the performance of a model in the context of SVD is the truncation error. This error arises due to the approximation of truncated SVD, which can be measured by calculating the difference between the original weight matrix $\mathbf{W}$ and the approximated matrix $\mathbf{W}_{k}$ as:
$\mathcal{L}_{\text{t}}=\|\mathbf{W}-\mathbf{W}_{k}\|_{F}.$ (2)
where $\|\|_{F}$ denotes the Frobenius norm. Exploiting the properties of SVD, we can further refine this error in terms of the singular values not included in the truncated matrix $\mathbf{W}_{k}$. Assuming the singular values in $\mathbf{\Sigma}$ are arranged in decreasing order, the truncation error can be described as the square root of the sum of the squares of the omitted singular values. This is also why truncated SVD is understood as the best low-rank approximation as it preserves the information most compactly [Eckart & Young, 1936]. Specifically, it is given by: $\mathcal{L}_{\text{t}}=\sum_{i=k+1}^{\min(m,n)}\delta_{i}^{2}$, where $\delta_{i}$ are the singular values that are not included in $\mathbf{\Sigma}_{k}$.
The choice of rank $k$ in SVD plays a crucial role in balancing the compression ratio and the resulting error. Specifically, opting for a smaller rank $k$ typically leads to a higher compression ratio but at the cost of increased truncation error and vice versa. However, the rate of this increase depends on the distribution of the singular values in the original matrix $\mathbf{W}$. If $\mathbf{W}$ possesses a few large singular values while the rest are considerably smaller, a low-rank approximation can effectively capture most of the matrix’s patterns. In such scenarios, even a relatively small $k$ can result in a modest truncation error. Conversely, in cases where the singular values of $\mathbf{W}$ do not exhibit rapid decay, achieving a low truncation error necessitates opting for a higher rank $k$. This variance in singular value distribution underscores the need for careful selection of $k$ to optimize both compression and model performance in the context of LLM decomposition.
3.2 Challenges of Compressing LLMs via SVD
The concept of decomposing the large matrices in LLMs (e.g., $4096\times 4096$ matrices are ubiquitous in LLaMA-7b [Touvron et al., 2023a]) into lower ranks presents a viable pathway for model compression. However, straightforward application of existing low-rank decomposition techniques, which typically necessitate training, proves ineffective for LLMs [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015, Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021]. The primary limitation of these methods is their reliance on training data and a large amount of optimization for rehabbing the decomposed models, contrasting with our objective to compress LLMs in a training-free manner. To address this gap, we conducted a comprehensive analysis to understand the failure points of traditional methods (i.e., SVD for large matrix in LLM) when applied to LLM compression. Our investigation revealed two primary challenges specific to post-training decomposition of LLMs.
Figure 2: Comparison between SVD and Activation-aware Singular Value Decomposition (ASVD). Outlier channels in input activations ($\mathbf{X}$) are highlight in red, and ASVD takes these into consideration, which can contribute to a reduction in output error.
Challenge 1: Sensitivity of Weight Variation: Recent studies in LLM post-training quantization [Lin et al., 2023, Wei et al., 2022, 2023, Frantar et al., 2022] have underscored the importance of considering activations when compressing LLM weights. This perspective shifts the focus from solely relying on the truncation error $\mathbf{L}_{\text{t}}$, which depends only on the model’s weights, to also accounting for the activations (see Eq.2). The rationale behind this is the critical role of outliers in activations within LLMs  [Lin et al., 2023, Wei et al., 2022, Kim et al., 2023]. Thus, for effective LLM decomposition, our objective optimization becomes:
$\mathbf{W}_{k}^{\star}$ $=\arg\min_{\mathbf{W}_{k}}\mathcal{L}(\mathbf{W}_{k}),$ (3)
$\arg\min_{\mathbf{W}_{k}}$ $\mathcal{L}(\mathbf{W}_{k})=\|\mathbf{W}_{k}\mathbf{X}-\mathbf{W}\mathbf{X}\|$
Here, $\mathbf{X}$ represents the input activations, which are cached from a small calibration set. This set is derived from the pre-training dataset to avoid overfitting to a specific task. Essentially, our objective is to ensure that the output of the decomposed LLM closely mimics the output of the original LLM, rather than merely aligning their weights. This approach prioritizes functional equivalence over structural similarity, recognizing that accurate output replication is more critical for maintaining the model’s post-decomposition performance. We define the variation in activations between the compressed matrix $\mathbf{W}_{k}$ and the original matrix $\mathbf{W}$ as:
$\Delta\mathbf{Y}=(\mathbf{W}_{k}-\mathbf{W})\mathbf{X}.$ (4)
To illustrate this concept, we visualize an example of $\mathbf{W}$, $\mathbf{W}_{k}$ (decomposed by simply SVD), $\mathbf{X}$, and the resulting variation in activations $\Delta\mathbf{Y}$ in Fig.2 (Top line). This visualization reveals a critical insight: even when the variation in weights $\Delta\mathbf{W}=\mathbf{W}-\mathbf{W}_{k}$ is relatively minor, the corresponding variation in activations $\Delta\mathbf{Y}$ can be huge. This significant variation in activations is a key factor why a straightforward SVD-based decomposition approach falls short in effectively decomposing LLMs. The activation variations, despite being derived from input activations large magnitude (not the weight variations), can lead to considerable changes in the whole model’s output, thereby undermining the decomposition’s efficacy.
Challenge 2: Singular Values Variations among Layers: The distribution of singular values within a matrix is indicative of its sparsity and, by extension, its sensitivity to certain types of information [Kim et al., 2015, 2019, Wen et al., 2017]. In LLMs, there is a notable variation in singular values across different layers. Specifically, some layers exhibit a concentration of large singular values, signifying less sensitivity to weight variation. This characteristic often correlates with these layers being easy to compress. Conversely, other layers in the LLMs display a more uniform distribution of smaller singular values. Such a pattern suggests a balanced contribution from various singular vector pairs. This variability in the distribution of singular values among layers presents a unique challenge, as it implies that each layer may require a tailored approach to decompose and maintain the overall functionality of the LLM.
These challenges underscore the necessity for innovative approaches specifically designed for the LLM decomposition. Our objective is to achieve efficient compression while circumventing the substantial computational and data demands associated with training-based methods. To address the first challenge, we introduce an Activation-aware SVD mechanism, which is detailed in Section 3.3. This method is designed to mitigate the impact of weight variation on activations. For the second challenge, we propose a Sensitivity-based Truncation Rank Searching mechanism, elaborated in Section 3.4, which adapts to the varying singular value distributions among different layers. These tailored strategies aim to ensure that the decompression maintains the functional integrity of the LLMs while optimizing the memory and computational efficiency.
3.3 Activation-aware Singular Value Decomposition
Our ASVD is a technique designed to refine the weight matrix $\mathbf{W}$ in LLMs by taking into account the influence of input activation channels. The process comprises the following three steps:
Scaling the Weight Matrix. The first step involves scaling the weight matrix $\mathbf{W}$ by a diagonal matrix $\mathbf{S}$. This matrix $\mathbf{S}$ is constructed to represent the impact of input channels on the weights, essentially adjusting $\mathbf{W}$ to better adapt with the activation patterns of the input $\mathbf{X}$, as demonstrated in Fig. 2 (Bottom line). The scaled weight matrix is denoted as $\mathbf{WS}$, and the equation is formulated as:
$\mathbf{W}=\mathbf{WSS}^{-1}=(\mathbf{WS})\mathbf{S}^{-1}.$ (5)
This scaling step is crucial as it modifies the weight matrix to reflect the varying significance of different input channels, setting the stage for a more effective and targeted decomposition process.
Applying SVD to the Scaled Matrix. After scaling the weight matrix, the next step is to apply SVD to the scaled matrix $\mathbf{WS}$. The SVD of $\mathbf{WS}$ is expressed as $\mathbf{WS}=\mathbf{U}^{\prime}\mathbf{\Sigma}^{\prime}{\mathbf{V}^{\prime}}^{T}$. To realize the matrix compression, we truncate this decomposition to retain only the first $k$ singular values and corresponding vectors. The truncated form of the decomposition is represented as:
$\mathbf{WS}\approx\mathbf{U}_{k}^{\prime}\mathbf{\Sigma}_{k}^{\prime}{\mathbf{V}_{k}^{\prime}}^{T}.$ (6)
This step is pivotal as it ensures that the most significant aspects of the scaled weight matrix are retained, while less critical information, which contributes minimally to the model’s output, is discarded.
Reconstructing the Approximated Weight Matrix. The final step in the ASVD process is to reconstruct an approximation of the original weight matrix. This reconstruction is accomplished by multiplying the truncated decomposition $\mathbf{U}_{k}^{\prime}\mathbf{\Sigma}_{k}^{\prime}{\mathbf{V}_{k}^{\prime}}^{T}$ with the inverse of the scaling matrix $\mathbf{S}^{-1}$. The multiplication of ${\mathbf{V}_{k}^{\prime}}^{T}$ with $\mathbf{S}^{-1}$ yields a new matrix ${\mathbf{V}_{k}^{\prime\prime}}^{T}$ (i.e., ${\mathbf{V}_{k}^{\prime\prime}}^{T}={\mathbf{V}_{k}^{\prime}}^{T}\mathbf{S}^{-1}$), which is the transformed version of the truncated right singular vectors considering the scaling effect. As a result, the final approximated weight matrix can be expressed as:
$\mathbf{W}\approx\mathbf{U}_{k}^{\prime}\mathbf{\Sigma}_{k}^{\prime}{\mathbf{V}_{k}^{\prime\prime}}^{T}.$ (7)
Our ASVD, with its focus on channel-wise activation outliers, strives to mitigate these outliers by appropriately scaling the corresponding weights. In this way, the activation variation can be mitigated as illustrated in Fig. 2. Note that the diagonal nature of $\mathbf{S}$ simplifies its inversion process, making $\mathbf{S}^{-1}$ computationally efficient. The subsequent challenge is to devise a method or criteria for accurately determining $\mathbf{S}$ so that it effectively aligns with the activation patterns observed in the LLM, ensuring that the scaling process optimally addresses the outlier activations.
Scaling Matrix $\mathbf{S}$ Exploration. Each diagonal element in the scaling matrix $\mathbf{S}$, particularly $\mathbf{S}_{ii}$ is pivotal in modulating the input activations as $(\mathbf{W}\mathbf{S})\mathbf{X}$. Each entry of $\mathbf{S}_{ii}$ specifically scales the influence of the $i$-th input channel. This scaling is instrumental in adjusting how each activation channel impacts the weight matrix during the decomposition process. The significance of $\mathbf{S}_{ii}$ is visualized in Fig.2 (bottom line), where its role in modulating channel influence is depicted. To effectively determine the values of $\mathbf{S}_{ii}$, various methodologies are explored. Each of these methods focuses on different characteristics of input activation, aiming to optimize $\mathbf{S}$ so that it accurately reflects the importance and influence of each input channel. This exploration is crucial for ensuring that the scaling aligns with the activation patterns observed in the LLM, thereby enhancing the effectiveness of the ASVD process.
Absolute Mean Value of Input Activation: This method calculates $\mathbf{S}_{ii}$ by calculating the absolute mean value of the activations in the $i$-th channel. Mathematically, if $\mathbf{X}_{i,:}$ represents the input activation values of the $i$-th channel, then $\mathbf{S}_{ii}$ is computed as:
$\mathbf{S}_{ii}=(\frac{1}{n}\sum_{j=1}^{n}|\mathbf{X}_{ij}|)^{\alpha},$ (8)
where $n$ is the total number of activations for the $i$-th channel and hyper-parameter $\alpha$ provides flexibility to adjust the level of activation sensitivity incorporated into the scaling. This method focuses on the average magnitude of activation in each channel, capturing the general intensity of activation signals irrespective of their positive or negative nature.
Absolute Maximum Value of Input Activation: Diverging from the mean-based approach, this method zeroes in on the peak activation within each channel. Under this approach, $\mathbf{S}_{ii}$ is determined by the highest absolute value observed in the activations of the $i$-th channel. The formal expression for this calculation is:
$\mathbf{S}_{ii}=(\max_{j}(|\mathbf{X}_{ij}|))^{\alpha}.$ (9)
This method places emphasis on the most significant, or peak, activation in each channel, positing it as a key indicator of the channel’s overall impact and importance. By focusing on the peak values, this approach underscores the channels with the most pronounced activation responses, considering them as crucial contributors to the network’s functionality.
3.4 Sensitivity-based Truncation Rank Searching
In addressing the second challenge - the variations of singular values among different layers of LLMs - we introduce the Sensitivity-based Truncation Rank Searching (STRS) method. This challenge arises from the fact that different layers in LLMs exhibit varying degrees of sensitivity to information compression, which is reflected in the distribution of their singular values. STRS aims to analyze and leverage this layer-specific sensitivity, particularly concerning how truncation of singular values impacts prediction accuracy.
Similar to any post-training compression methods [Dettmers et al., 2022, Frantar et al., 2022, Frantar & Alistarh, 2023], we first collect a small but representative calibration dataset. This dataset is instrumental in evaluating the layers’ responsiveness to different truncation ranks. For the calibration set, we select a sample of sequences from a source such as Wikitext [Merity et al., 2016]. In the realm of NLP, perplexity is a key metric for assessing how effectively a language model predicts a sequence of tokens [Brown et al., 2020]. Within the scope of our analysis, “sensitivity” refers to the reduction in perplexity observed in the calibration dataset following the decomposition of the model.
The core of the sensitivity evaluation process involves an in-depth exploration of how the neural network reacts to varying levels of truncation. We define a set of potential truncation ratios, denoted as $R$, which include values like $\{0.1,0.2,\cdots,0.9\}$. These ratios determine the fraction of the rank $k$ preserved during the SVD truncation for a weight matrix with dimensions $m\times n$. The truncation ratio $r$ is mathematically defined as:
$r=\frac{km+kn}{mn},r\in R.$ (10)
For each linear layer in the LLM, we iterate through these candidate ratios. At each ratio, truncated SVD is applied to the layer’s weight matrix, temporarily replacing the original layer in the model with this truncated version. The model’s perplexity is then evaluated on the calibration dataset with these adjustments. This iterative process allows us to determine the truncation ratio that optimally balances model compression with minimal impact on prediction accuracy.
Figure 3: Perplexity across Various Linear Layers and Parameter Ratios on LLaMA-2-7b.
This detailed exploration of sensitivity across various truncation levels provides essential insights into each layer’s performance dynamics, informing the optimization and decision-making processes in model compression. As illustrated in Fig. 3, there are noticeable variations in sensitivity among the different layers of the network. Three key observations emerge from this analysis: (i) Inversely Proportional Relationship: We observe that perplexity is inversely proportional to the parameter ratio. Specifically, lower parameter ratios tend to result in higher perplexity scores. This finding highlights the critical trade-off between model compression and reasoning performance, signaling the importance of careful parameter selection to maintain a balance. (ii) Higher Sensitivity in MLP Layers: Layers within Multi-Layer Perceptrons (MLP) [Vaswani et al., 2017] demonstrate higher sensitivity compared to those in feed forward modules. This distinction is crucial in guiding the compression strategy, indicating where more cautious truncation is necessary. (iii) Variable Sensitivity Among Layers: Some layers exhibit relatively lower sensitivity, indicating potential for more aggressive compression without significant performance degradation. This variation suggests the presence of redundancy within LLMs, presenting opportunities for more efficient compression approaches. Targeting these less-sensitive layers could lead to models that are both compact and computationally efficient.
Post sensitivity evaluation, we compile a list of tuples (layer, truncation rank, sensitivity) and sort it according to sensitivity. To determine an optimal rank configuration that aligns with these insights, we propose a binary search algorithm. The details of this algorithm are outlined in Algorithm 1, designed to navigate the complexity of selecting the most effective truncation ranks for each layer, based on their sensitivity and contribution to the overall model performance. More detailed explanations about this algorithm can be found in the Appendix.
3.5 Absorbing Singular Values
After we decompose a matrix via ASVD, we can represent the weight matrix as a product of three matrices, i.e., $\mathbf{W}\approx\mathbf{U}_{k}\mathbf{\Sigma}_{k}\mathbf{V}_{k}^{T}$. Thanks to the diagonal nature of matrix $\mathbf{\Sigma}_{k}$, we can further optimize the inference process. Specifically, we can efficiently absorb the singular values in $\mathbf{\Sigma}_{k}$ into the matrices $U_{k}$ and $V_{k}^{T}$. We achieve this fusion using the following strategy: $\mathbf{A}_{k}=\mathbf{U}_{k}\sqrt{\mathbf{\Sigma}_{k}}$ and $\mathbf{B}_{k}=\sqrt{\mathbf{\Sigma}_{k}}\mathbf{V}_{k}^{T}$. Consequently, we obtain a more computationally efficient matrix operation  [^1] See Appendix for details.:
$\mathbf{Y}=\mathbf{W}\mathbf{X}\approx\mathbf{A}_{k}(\mathbf{B}_{k}\mathbf{X})$ (11)
Compared to the methods of fusing the singular values $\mathbf{\Sigma}_{k}$ solely into either $\mathbf{U}$ or $\mathbf{V}$ matrices, our proposed fusion technique offers significant advantages in terms of weight quantization, as demonstrated in Table 1. Our approach involves evenly distributing the singular values from the diagonal matrix $\mathbf{\Sigma}_{k}$ into both $\mathbf{U}_{k}$ and $\mathbf{V}_{k}^{T}$ matrices. This ensures a more uniform distribution of $\mathbf{A}_{k}$ and $\mathbf{B}_{k}$, leading to a reduction in the disparity across different channels and reducing the quantization error.
Table 1: Perplexity on Wikitext-2 under different absorbing strategies after ASVD on OPT-125m.
param ratio weight quant absorbed by UV absorbed by U absorbed by V
0.9 INT6 37.58 39.67 40.62
0.85 INT6 60.44 64.19 61.02
4 Experiments
In this section, we assess the effectiveness of ASVD by conducting experiments on LLaMA [Touvron et al., 2023a] and LLaMA-2 [Touvron et al., 2023b], and presenting results on various tasks, such as Perplexity in WIKI [Merity et al., 2016] and MMLU [Hendrycks et al., 2020].
4.1 Hyper-parameters Exploration
In our study, we initiated an exploration of hyper-parameters in ASVD, focusing on the activation channel significance metric and the control factor $\alpha$. This exploration were conducted on OPT-125m, a relatively small network that facilitates rapid evaluation.
We rigorously explored the control factor $\alpha$ at various settings: 0.1, 0.25, 0.5, 1, and 2. This exploration aimed to understand how varying $\alpha$ influences the performance and parameter efficiency of the network. Additionally, we investigated two methods for quantifying activation significance: Absolute Mean Value of Input Activation and Absolute Maximum Value of Input Activation. These methods are crucial in determining the most effective approach for activation channel significance evaluation. We set a target parameters ratio of 0.9. Utilizing the binary search approach for truncation ranks, we report the perplexity on Wikitext2 test set after compression.The results of our experiments are summarized in Table 2.
From the data presented in the table, we observe that both activation-aware methods show superior performance compared to standard SVD+STRS. We also notice that Lower and higher values of $\alpha$ (0.1 and 2) exhibit lower performance, while mid-range values (0.5) lead to better performance, and the Absolute Mean Value method consistently outperforms the Absolute Max Value method. Therefore, based on our observations, we chose $\alpha=0.5$ and the Absolute Mean Value method for setting the scaling matrix $S$ in the ASVD process in the following experiments.
4.2 Evaluation of ASVD on LLaMA Models
We conducted a comprehensive evaluation of Adaptive Singular Value Decomposition (ASVD) on two LLaMA models: LLaMA-7b and LLaMA-2-7b, each comprising 7 billion parameters. For each model, we selected 32 calibration datasets, each containing 2048 tokens, from the Wikitext dataset to assess the layer-wise sensitivity.
Table 2: Perplexity on Wikitext2 for exploring hyper-parameters on OPT-125m.
$\alpha$ 0.1 0.25 0.5 1 2
SVD+STRS 103.39
ASVD abs mean 47.54 37.12 36.89 41.53 43.81
ASVD abs max 52.63 47.17 40.14 41.94 52.55
Figure 4: Perplexity trends following ASVD compression on LLaMA-2-13b, LLaMA-2-7b and LLaMA-7b.
Table 3: Performance under various compression scenarios. Param ratio indicates the proportion of parameters remaining after decomposition. MMLU results are 0-shot.
method param ratio MMLU wiki ptb
LLaMA-7b
original 1 30.76% 5.68 29.63
SVD 0.95 22.98% 2800.94 5458.67
SVD+STRS 0.95 23.92% 136.05 183.92
SVD+STRS 0.9 23.54% 698.66 262.03
ASVD 0.95 30.26% 5.78 32.64
ASVD 0.9 29.67% 6.09 37.80
ASVD 0.85 29.70% 6.80 52.11
ASVD 0.8 27.85% 8.89 88.09
ASVD 0.75 24.94% 14.51 212.80
LLaMA-2-7b
original 1 34.86% 5.47 20.82
SVD 0.95 - nan nan
SVD+STRS 0.95 24.78% 46.79 363.37
SVD+STRS 0.9 24.31% 114.45 27660.43
ASVD 0.95 33.24% 5.64 23.98
ASVD 0.9 32.58% 5.93 32.63
ASVD 0.85 31.57% 6.74 59.84
ASVD 0.8 28.15% 8.91 114.70
ASVD 0.75 25.97% 18.97 432.57
LLaMA-2-13b
original 1 40.16% 4.88 29.21
SVD 0.95 - nan nan
SVD+STRS 0.95 24.86% 167.63 567.02
SVD+STRS 0.9 - nan nan
ASVD 0.95 39.52% 4.94 31.93
ASVD 0.9 40.04% 5.12 34.03
ASVD 0.85 37.95% 5.54 39.32
ASVD 0.8 34.63% 6.53 59.68
ASVD 0.75 28.59% 8.71 110.10
Our methodology involved setting varying thresholds binary searching process, enabling us to observe the impact of different compression levels on model performance. This approach resulted in a range of compressed networks, each characterized by a unique compression ratio.
We evaluated the performance of these compressed networks using perplexity as the primary metric, focusing on two datasets: Wikitext-2 and the Penn Treebank (PTB). The objective was to understand the relationship between compression ratio and model performance. The results, illustrated in Figure 4, reveal several key insights: (1) As the parameter ratio decreases, there is a corresponding increase in perplexity. As more parameters are compressed, the more error is introduced by SVD. This reduces representational capacity of the network, thereby resulting in higher perplexity scores. (2) A plateau region is observed when the parameter ratio exceeds 0.9. In this range, ASVD predominantly decompresses the less sensitive layers, resulting in minimal impact on prediction accuracy. (3) Below a parameter ratio of 0.85, there is a rapid increase in perplexity, indicating that the more sensitive layers are being decompressed to a lower truncation rank, adversely affecting the model’s performance.
We present a detailed analysis of the performance of compressed networks at various parameter ratios, as summarized in Table 3. This analysis aims to elucidate the impact of different compression techniques on network efficiency and accuracy.
Table 3 displays the performance metrics for two LLaMA models, LLaMA-7b and LLaMA-2-7b, under several compression scenarios. These metrics include Massive Multitask Language Understanding (MMLU) zero-shot evaluation, perplexity on the Wikitext dataset (wiki), and perplexity on the Penn Treebank dataset (ptb).
Our observations indicate significant variations in performance as a function of the parameter ratio and the compression method employed. Specifically, the table highlights the performance of each model when using standard Singular Value Decomposition (SVD), SVD with binary search for truncation ranks (SVD*), and ASVD at different parameter ratios ranging from 0.75 to 0.95.
We also compare ASVD with FWSVD [Hsu et al., 2022], which uses Fisher information to weigh the importance of parameters affecting the model prediction. Note that FWSVD is training-required. As shown in Table 4, our method can outperform FWSVD comprehensively.
Table 4: Comparing with FWSVD on LLaMA-7b. FWSVD* denotes Fisher information weighted SVD.
param ratio 0.95 0.9 0.85 0.8
LLaMA-7b
FWSVD+STRS wiki 5.86 6.32 7.48 10.70
ASVD 5.78 6.09 6.80 8.89
FWSVD+STRS ptb 34.33 38.05 58.75 125.80
ASVD 32.64 37.80 52.11 88.09
LLaMA-2-7b
FWSVD+STRS wiki 5.59 6.12 8.01 13.07
ASVD 5.64 5.93 6.74 8.91
FWSVD+STRS ptb 25.06 36.58 105.53 222.03
ASVD 23.98 32.63 59.84 114.70
4.3 Integrating ASVD with Quantization
Table 5: Combining weight quantization with ASVD. Param ratio indicates the proportion of parameters remaining after ASVD, with 1 implying no decomposition.
param ratio dtype LLaMA-2-7b LLaMA-2-13b
wiki ptb wiki ptb
1 FP16 5.47 20.82 4.88 29.15
1 INT8 5.48 20.82 4.88 29.12
1 INT6 5.52 20.97 4.90 29.03
1 NF4 5.65 22.70 4.98 30.31
0.95 FP16 5.64 23.98 4.94 31.93
0.95 INT8 5.64 23.95 4.95 31.67
0.95 INT6 5.70 25.10 4.97 31.97
0.95 NF4 5.83 35.91 5.08 33.89
0.9 FP16 5.93 32.63 5.12 34.03
0.9 INT8 5.94 32.19 5.11 33.64
0.9 INT6 6.02 32.83 5.16 33.66
0.9 NF4 6.20 40.82 5.31 34.93
0.85 FP16 6.74 59.84 5.54 39.32
0.85 INT8 6.73 63.76 5.56 40.02
0.85 INT6 6.90 66.88 5.62 38.15
0.85 NF4 7.43 427.59 5.90 44.49
This section explores the integration of ASVD with quantization techniques to compress large language models. While our primary focus is on ASVD, we investigate the synergy between ASVD and simple quantization methods, Round-To-Nearest (RTN) and 4-bit NormalFloat (NF4) [Dettmers et al., 2023], as a preliminary step. Future work could extend this investigation to more advanced quantization methods.
Our experimental framework involves two stages. Firstly, we apply ASVD to decompose the network. Subsequently, we quantize the decomposed weights. We employ per-channel asymmetric quantization, targeting two quantization levels: 8-bit and 6-bit. We employ NF4 for quantizing to 4-bit. The impact of these quantization levels on network performance is then assessed.
Table 5 summarizes the results of our experiments on LLaMA-7b and LLaMA-2-7b models. The following observations were made: 8-bit Quantization: The results indicate that 8-bit quantization has a negligible impact on model performance, both for the original and the ASVD-compressed networks. 6-bit Quantization: A slight degradation in prediction accuracy is observed with 6-bit quantization. This effect is more pronounced in networks with higher compression (lower parameter ratios), suggesting a compounded impact of compression and lower bit quantization. 4-bit Quantization: Upon quantizing the network into NF4, a further deterioration in prediction accuracy is observed. When param ratio is greater than 0.9, the performance decline attributed to quantization is approximately consistent with that of the non-decomposed network. For instance, in the case of LLaMA-2-7b without decomposition, the NF4 quantization results in a wiki PPL decrease of 0.18 compared to FP16. Similarly, at a param ratio of 0.95, the NF4 quantization yields a PPL decrease of 0.19 compared to FP16. In summary, the findings suggest that ASVD is compatible with weight quantization techniques.
4.4 Decomposed Network Analysis
Figure 5: Per-type parameters ratio and per-block parameters ratio on LLaMA-2-7b after ASVD compression.
In this section, we conduct a detailed analysis of the decomposed network. Figure 5 presents the per-type parameters ratio and per-block parameters ratio. Observing the plot, we note that parameters in the MLP components (gate projection, up projection, and down projection) exhibit minimal compression. In MHA, the V projection layer experiences relatively small compression, whereas q projection and k projection can be significantly compressed, indicating redundancy in these components. Turning our attention to the per-block compression ratio, we find that the first layer can undergo substantial compression. In contrast, the compression ratios for the other layers, except for two middle layers, show similar compression rates.
5 Conclusion
This study introduces a training-free approach to compress the Large Language Models (LLMs), enhancing their usability across diverse computing environments. The Activation-aware Singular Value Decomposition (ASVD) and Sensitivity-based Truncation Rank Searching (STRS), central to our approach, adeptly addresses activation outliers and varying layer sensitivities, leading to more accurate and efficient decomposition. Our experiments validate that ASVD can compress LLMs by 10%-20% without compromising their reasoning capabilities. Furthermore, ASVD’s compatibility with other LLM compression methods underscores its potential as a versatile tool for LLM deployment.
Appendix A Appendix
Algorithm 1 Binary Search for Truncation Ranks (performance target)
  Input: List of tuples (layer, truncation rank, sensitivity), perplexity target and calibration dataset
  Output: Optimal rank configuration for each layer
  Sort the list by sensitivity in ascending order
  Initialize pointers: $p_{L}=0$, $p_{H}=\text{length of list}-1$
  $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$
  while $p_{L}\neq p_{H}$ do
     for each layer in the list do
        Initialize $r=\infty$
        for each tuple in the list starting from $p_{M}$ to the end do
           if tuple’s layer is the same as the current layer then
[…]
           end if
        end for
        if $r=\infty$ then
           Not decompose the layer
        else
           Decompose the layer with truncation rank $r$
        end if
     end for
     Evaluate network performance (perplexity) on calibration dataset
     if perplexity $\leq$ perplexity target then
        $p_{H}=p_{M}$
     else
        $p_{L}=p_{M}+1$
     end if
     Update $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$
  end while
A.1 Inference Cost with Decomposed LLMs
Regarding the computational aspect, let’s consider the input matrix $\mathbf{X}\in\mathbb{R}^{n\times t}$ and the weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$. In the original linear layer, the matrix multiplication is represented as $\mathbf{Y}=\mathbf{X}\mathbf{W}$. The number of Multiply-Accumulate (MAC) operations, denoted as $C$, in the original linear layer can be computed as: $C=tmn$. After the ASVD decomposition, the matrix multiplication transforms into $\mathbf{Y}=\mathbf{A}_{k}\mathbf{B}_{k}$. To analyze the computational efficiency, we calculate the MAC operations, denoted as $C_{k}$, for this decomposed form. The computation for $C_{k}$ is given by: $C_{k}=tkm+tkn$
This computation ratio can be expressed as the ratio of $C_{k}$ to $C$, which is equivalent to the parameter ratio:
$\frac{C_{k}}{C}=\frac{km+kn}{nm}$ (12)
Remarkably, this computation ratio mirrors the weight number compression ratio, highlighting the efficient use of computational resources achieved through ASVD. In summary, ASVD can not only reduce the weight storage and weight transferring overheads in LLM deployment but also reduce the computation required by LLM inference.
A.2 Binary Search for Truncation Ranks
We have the option to employ either a performance target or parameters target for our search. In the case of a performance target, our objective is to identify the truncation rank configuration that ensures the compressed network attains the desired performance, such as achieving a specific perplexity. Alternatively, in the pursuit of a parameters target, our goal is to identify the truncation ranks that result in the network attaining the specified target parameters.
The algorithm of performance target is shown in Algorithm 1. Initially, the low pointer ($p_{L}$) is positioned at the start of the list, while the high pointer ($p_{H}$) is set at the list’s end. The middle pointer ($p_{M}$), as the name suggests, is placed midway between $p_{L}$ and $p_{H}$, calculated as $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$. During each iteration of the binary search, we adjust the truncation rank for each layer. Specifically, for a given layer, its truncation rank is set to the smallest rank found to the right of the middle pointer ($p_{M}$) in our list.
Following this adjustment, we evaluate the network’s performance using the updated configuration on a calibration dataset. The primary metric for assessment is perplexity. Should the perplexity fall within or below a pre-established threshold, we move the high pointer ($p_{H}$) to the middle position ($p_{M}$). This action indicates our search for a configuration with a potentially lower rank that still adheres to performance standards. Conversely, if the perplexity exceeds our maximum acceptable threshold, we shift the low pointer ($p_{L}$) to ($p_{M}+1$). This adjustment signifies the need to increase the truncation ranks to maintain or enhance performance levels. The binary searching will converge to an optimal configuration of truncation ranks for each layer that balances compression ratio and perplexity.
Algorithm 2 Binary Search for Truncation Ranks (parameters target)
  Input: List of tuples (layer, truncation rank, sensitivity) and parameters target
  Output: Optimal truncation rank configuration for each layer
  Sort the list by sensitivity in ascending order
  Initialize pointers: $p_{L}=0$, $p_{H}=\text{length of list}-1$
  $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$
  while $p_{L}\neq p_{H}$ do
     for each layer in the list do
        Initialize $r=\infty$
        for each tuple in the list starting from $p_{M}$ to the end do
           if tuple’s layer is the same as the current layer then
[…]
           end if
        end for
        if $r=\infty$ then
           Do not modify the truncation rank for the layer
        else
           Set the truncation rank for the layer to $r$
        end if
     end for
     Calculate the parameters after compression
     if parameters $\leq$ parameters target then
        $p_{H}=p_{M}$
     else
        $p_{L}=p_{M}+1$
     end if
     Update $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$
  end while
The algorithm of parameters target is shown in Algorithm 2. It doesn’t need calibration dataset.